Change the following point from polar to rectangular coordinate.

(3/2, n/12)

II. Evaluate f²(3ײ- 6x ‐2)dx



1
Expert's answer
2022-01-17T13:28:41-0500

a) "x=r\cos \theta, y=r\sin \theta"


"x=\dfrac{3}{2}\cos\dfrac{\pi}{12}=\dfrac{3}{2}(\dfrac{\sqrt{6}+\sqrt{2}}{4})=\dfrac{3(\sqrt{6}+\sqrt{2})}{8}"

"y=\dfrac{3}{2}\sin\dfrac{\pi}{12}=\dfrac{3}{2}(\dfrac{\sqrt{6}-\sqrt{2}}{4})=\dfrac{3(\sqrt{6}-\sqrt{2})}{8}"

"\big(\dfrac{3(\sqrt{6}+\sqrt{2})}{8}, \dfrac{3(\sqrt{6}-\sqrt{2})}{8}\big)"

b)

"\int (3x^2-6x-2)dx=x^3-3x^2-2x+C"